# Thermal Endurance of Suspended Thin-Film Lithium Niobate up to 800 C

### v1

@@ 4.2 Device Performance
Overall, the resonant characteristics follow the same trend as earlier annealing rounds, with the exception of a resonant frequency downshift, and the Q experiences significant improvements in the late annealing rounds. The wideband admittance of the example resonator is shown in Fig. 14. The admittances of the individual modes are shown in Figs. 15-17, and the extracted resonance parameters are given in Tables IX-XI. After the 600 C anneal, the resonant frequency is lower compared to the post-550 C anneal round resonant frequency. This is the same annealing round where the Ti disappears from the DAD rocking cruves. In modes 2 and 3, the resonant frequency increases after a 650 C anneal and subsequent anneal rounds, but mode 1’s resonant frequency only starts increasing after the 700 C anneal. The overall increase in resonant frequency across the modes can be attributed to increased stress and strain after annealing. After the 700 C anneal, all modes exhibit Qs larger than their initial value. Modes 2 and 3 have significant Q improvements of over 100% compared to their original value after the 700 C anneal, as well as after the 650 C anneal for mode 2. In addition to the metal resistivity decreasing for functional devices, some of the resonator’s anchors break (Fig. 10). Less electrical energy is lost because of lower resistivity, and less mechanical energy is lost because the resonator can vibrate more freely compared to when all the anchors are intact. The origins of the frequency drifting and Q changes are quite intriguing for future work, but are out of the scope for this work on thermal endurance.
@@
After the 800 C anneal, the example resonator no longer functions. Although in Fig. 14 a small resonance can be seen around mode 1’s frequency, the much lower capacitance indicates broken anchors because only part of the signal gets capacitively fed into the resonator via the cracked point; furthermore, modes 2 and 3 are no longer visible and the admittance drops significantly compared to previous annealing rounds. Out of 60 measured resonators, 33 (55%) became inoperable after the 700 C anneal (Table XII). This suggests the Si cracking, which also occurs after the 700 C anneal, is a major failure point of the devices. Prior to the 700 C anneal, the only devices that fail are anchored devices. Additionally, 20 resonators had anchors, while 40 resonators had no anchors. Although some resonators are able to operate after a 750 C anneal, it is not near LN’s Curie temperature of 1200 C. Improvements can be made in the anchor design for better stability, in the metal design to delay the metal failure, and different materials can be considered to better match the CTE of LN. Nevertheless, functionality after a 750 C anneal demonstrates the suspended thin-film LN platform’s viability for high-temperature applications.
@@ 5 Conclusion
This paper reports a suspended thin-film LN acoustic resonator among 60 in a high-temperature study surviving annealing temperatures of up to 750 C, marking the highest temperature point the platform has been demonstrated in. Despite thermal damage, including cracking, loss of conductivity, and lower LN quality, the survival of the resonator demonstrates the strong thermal endurance of the platform. The platform can be used for MEMS in harsh thermal environments and can push other suspended LN systems, such as optical systems, into higher temperature applications.

### v2

@@ 4.2 Device Performance
Overall, the resonant characteristics follow the same trend as earlier annealing rounds, with the exception of a resonant frequency downshift, and the Q experiences significant improvements in the late annealing rounds. The wideband admittance of the example resonator is shown in Fig. 14. The admittances of the individual modes are shown in Figs. 15-17, and the extracted resonance parameters are given in Tables XI-XIII. After the 600 C anneal, the resonant frequency is lower compared to the post-550 C anneal round resonant frequency. This is the same annealing round where the Ti disappears from the DAD rocking cruves. In modes 2 and 3, the resonant frequency increases after a 650 C anneal and subsequent anneal rounds, but mode 1’s resonant frequency only starts increasing after the 700 C anneal. The overall increase in resonant frequency across the modes can be attributed to increased stress and strain after annealing, which increases the speed of sound in the LN. Since the acoustic wavelength remains fixed due to the interdigitated electrode dimensions, the resonant frequency increases. After the 700 C anneal, all modes exhibit Qs larger than their initial value. Modes 2 and 3 have significant Q improvements of over 100% compared to their original value after the 700 C anneal, as well as after the 650 C anneal for mode 2. In addition to the metal resistivity decreasing for functional devices, some of the resonator’s anchors break (Fig. 10). Less electrical energy is lost because of lower resistivity, and less mechanical energy is lost because the resonator can vibrate more freely compared to when all the anchors are intact. Overall, the changes in Q do not directly correlate with the change in metal resistivity, suggesting a significant Q contribution from the LN’s elasticity as it changes after annealing rounds. The origins of the frequency drifting and Q changes are quite intriguing for future work, but are out of the scope for this work on thermal endurance. Future work on quantifying high-temperature damage through multi-physics coupling models could be of interest.
After the 800 C anneal, the example resonator no longer functions. Although in Fig. 14 a small resonance can be seen around mode 1’s frequency, the much lower capacitance indicates broken anchors because only part of the signal gets capacitively fed into the resonator via the cracked point; furthermore, modes 2 and 3 are no longer visible, and the admittance drops significantly compared to previous annealing rounds. Out of 60 measured resonators, 33 (55%) became inoperable after the 700 C anneal (Table XIV). This suggests the Si cracking, which also occurs after the 700 C anneal, is a major failure point of the devices. Prior to the 700 C anneal, the only devices that fail are anchored devices. Additionally, 20 resonators had anchors, while 40 resonators had no anchors. Although some resonators are able to operate after a 750 C anneal, it is not near LN’s Curie temperature of 1200 C. The device performance gradually degrades at higher temperatures due to a combination of electrode and piezoelectric layer degradation, and also structural damage due to thermal mismatch. Improvements can be made to the anchor design for enhanced stability, the metal design to delay metal failure, and different materials can be considered to better match the CTE of LN. Future studies can focus on identifying stack and device configurations for better thermal endurance. Future broader works on higher temperature systems, e.g., via developing higher temperature piezoelectric and electrode materials, would be of research interest, as well as advanced characterization and modeling methods. Nevertheless, functionality after a 750 C anneal demonstrates the suspended thin-film LN platform’s viability for high-temperature applications.
@@
Figure 18: Distributions of percent change in resonant frequency shift relative to the 250 °C anneal round for S0 modes near 500 MHz. The median of each distribution is depicted with a circle. The frequency shifts are normalized to the 250 °C anneal round due to the large variance of frequency shifts after the first anneal. The frequencies of subsequent anneals stay within one percent of the resonant frequency after the initial anneal. After the 550 °C and following anneal rounds, there are fewer points in the distributions due to device failure.
@@ 5 Conclusion
This paper quantified the thermal endurance of suspended thin-film X-cut stoichiometric LN by characterizing Lamb wave S0 mode acoustic resonators and meandering line resistivity structures with a Pt/Ti metal stack all on the same substrate. The devices were subjected to incremental annealing rounds from 250 C to 800 C. The resonators experience an overall increase in resonant frequency, maintaining a frequency within 1% of their initial anneal value. After the 600 C anneal, the Ti disappears from XRD measurements and a LiNb3O8 phase forms. After the 650 C anneal, metal discontinuity from hole formation was observed. After the 700 C anneal, many devices failed due to LN cracking and delamination. Despite these damages, the survival of some resonators after a 750 C anneal demonstrates the platform’s strong thermal endurance. These results provide baseline data to guide robust device and stack choices for MEMS and related LN systems operating in harsh thermal environments.